A tornado in Kansas has a radius of 1000 m and the pressure inside the core of the storm is about 15 kPa lower than the ambient atmosphere pressurei.e., it is only about 85 kPa compared to a typical 100 kPa). Assuming these values -3 this cyclostrophic motion.
Answer #1

Given that,

R = 1000 m

dp / dn = -15 kPa

\alpha = -1 / density = 1 / 1.2

\alpha = 0.833

Fomula for cyclotrophic wind,

v^2 / R = -\alpha * dp / dn

v^2 / 1000 = -0.833 * (-15000)

v^2= 12500000

v = 3535.53 m/s

Hence, typical wind velocity = 3535.53 m/s